Frequency of yellow light emitted by sodium  lamp  is 5.17×10^14Hz.  The wave  number  of yellow light is? 
 
 

Grade:12th pass

1 Answers

Khimraj
5 years ago
wave number =1/wavelength = frequency/velocity 
= 5.17*1014/3*108 = 1.72*106
Hope it clears.................................